OpenStudy (anonymous):

Solve by the addition method. (You may first have to multiply both sides of one equation by -1) Problem: 3x+y=13 x+y=3

OpenStudy (anonymous):

mult both sides of 2nd equation with -1 then add 2 equations together, so y is gone, now u can find x, then plug result found for x in any of the 2 equations to get y.

OpenStudy (anonymous):

so would I do it like this: x+ (-1) y=3(-1) ? then solve

OpenStudy (anonymous):

you need x to times -1 too, like this (-1)(x+y)=3(-1) (-1)x + (-1)y = 3(-1)

OpenStudy (anonymous):

so x = -5?

OpenStudy (anonymous):

nope, x=5

OpenStudy (anonymous):

How would you do this equation: x + 2y = -2 4x + 2y = -17 ?

OpenStudy (anonymous):

would x be -5 and y be -3/2?

OpenStudy (anonymous):

you get the last one right? So its the same for this one. time the first equation with (-1) (-1)(x+2y)=(-2)(-1) -x -2y = 2 then add this to the 2nd

OpenStudy (anonymous):

you got x right but y is wrong

OpenStudy (anonymous):

check ur work, y should be 3/2

OpenStudy (anonymous):

I accidently put the - I had 3/2. See if you can help me with this problem: -5=10x+3y -1= 2x-3y?

OpenStudy (anonymous):

lol this is even simpler, you dont need to change anything, just add them straight

myininaya (myininaya):

x=5

myininaya (myininaya):

not -5

OpenStudy (anonymous):

I told you the problem wrong, I'm sorry : -5=-10x+3y 1=2x-3y

myininaya (myininaya):

its still a good problem

OpenStudy (anonymous):

same thing -5 = -10x + 3y + 1 = 2x + (-3y) __________________________________ -4 = -8x + 0 then x=2, so find y

OpenStudy (anonymous):

lol sorry 1/2, I should go back to kindergarten

myininaya (myininaya):

lol i was just fixing to say that

myininaya (myininaya):

not the kindergarten part

OpenStudy (anonymous):

yep, that was stupid

OpenStudy (anonymous):

So is y 5?

OpenStudy (anonymous):

which question do you mean? the last one with -5=-10x...., y should be 0